# DeepPolar: Inventing Nonlinear Large-Kernel Polar Codes via Deep Learning

### v1

@@ 4.2 Results
DeepPolar codes outperform Polar codes. As highlighted in Fig. 5, DeepPolar codes achieve enhanced performance over traditional Polar codes with SC decoding and Reed-Muller codes with Dumer decoding across a broad range of SNRs in the presence of additive white Gaussian noise (AWGN), in terms of bit error rate (BER). Notably, DeepPolar (256,37,$\ell=16$) outperforms KO(8,2) (Fig. 5a), the state-of-the-art neural code mapping $37$ information bits to a length-$256$ codeword (Makkuva et al., 2021) in the short-to-medium blocklength regime. DeepPolar codes, owing to their Polar-like encoding framework, accommodate a much wider range of rates than KO codes that rely on the algebraic structure of RM codes. This versatility is illustrated in Fig. 5b and Fig. 5c, where DeepPolar codes outperform Polar codes across diverse rates and SNRs (cf. KO codes and RM codes do not exist for these rates). Further, DeepPolar is robust to non-AWGN deviations, achieving gains over Polar codes on fading channels and radar noise (App. D)
@@
Effect of kernel size. Fig. 6 highlights that scaling the kernel size $\ell$ is crucial for effective training and DeepPolar outperforming baselines. An increase in kernel size expands the function space the encoder can represent, facilitating learning more robust representations. However, empirical experiments reveal that using $\ell=\sqrt{n}$ , which results in an encoding and decoding tree of depth 2, is a good heuristic. Indeed, Fig. 1 highlights significant improvements in performance upto $\ell=16,32$, whereas further scaling $\ell=64$ leads to a decline. This trend indicates a bias-variance tradeoff, suggesting the ideal kernel size strikes a balance between model complexity and its generalization capabilities.
@@ 4.3 Interpretation
Figure 7: Unlike Polar codes, the distribution of pairwise distribution between codewords of DeepPolar shows a strong resemblance to the Gaussian codebook. DeepPolar-binary code retains a Gaussian-like distance profile. (pdfs and pmfs have been renormalized for better visualization.)
@@
Interpreting the decoder. In our study, the training loss used is binary cross entropy (BCE), which serves as a surrogate for the bit error rate (BER). However, it is important to recognize that optimizing for BER does not necessarily translate to improved block error rate (BLER), an important figure of merit in practical systems. Our observations, as indicated in Fig. 8a, show that DeepPolar underperforms in BLER compared to baseline methods (See App. F for more results). The BLER, defined as $\mathbb{P}[\hat{u}\neq u]$, can be expressed as the cumulative probability of bit errors occurring at each position when no errors were made in the previous ones:
$\mathbb{P}[\hat{u}\neq u]=\sum_{i=1}^{k}\mathbb{P}[\hat{m}_{i}\neq m_{i}|\hat{m}_{1:i-1}=m_{1:i-1}]$ (3)
An analysis of error distribution (Fig. 8b) reveals notable differences between SC and DeepPolar-SC decoders despite both employing a sequential decoding approach. Due to the effect of channel polarization, most errors in the polar decoding tend to occur at bit position 0, which predominantly drives the BLER. In contrast, errors in DeepPolar decoding are more evenly distributed across various bit positions. This uniformity is a consequence of the BCE loss prioritizing BER over BLER. This trade-off leads to a marginal decrease in BLER performance. Identifying a stable loss function directly targeting BLER, is an interesting direction for future research.
@@ 4.4 Ablation studies
DeepPolar-Binary Compared to canonical coding schemes, a real-valued coding scheme like DeepPolar offers distinct advantages by integrating modulation and coding schemes. However, practical systems may have hard symbol-level power constraints, necessitating code bits to be integers $x_{i}\in\{-1,1\}$. Hence, it is beneficial to have the ability to learn a binary code while maintaining the structure of DeepPolar.
To achieve this, a trained DeepPolar model is fine-tuned with a Straight Through Estimator (STE) (Hubara et al., 2016) and a binarization module, resulting in a binary version, DeepPolar-binary, akin to the methodology in (Jiang et al., 2019b). Fig. 5a highlights that there is a loss incurred with respect to DeepPolar. This underscores the contribution of joint coding and modulation to DeepPolar’s performance. Nevertheless, DeepPolar-binary not only surpasses the performance of the traditional Polar code but also closely matches that of KO codes. Additionally, the distance profile of DeepPolar-binary (Fig. 7) indicates it preserves Gaussian codebook-like distance properties, reinforcing the potential of non-linear binary codes derived from polarization-inducing structures. This exploration opens avenues for future research, particularly in interpreting these codes to develop novel, efficient encoders and decoders leveraging non-linear polarization-based methods.
@@ 4.5 Computational Complexity
In our study, we introduce a novel non-linear generalization of the Polar code’s encoding and decoding structures and achieve substantial improvements in reliability. Another essential aspect of evaluating an algorithm is the computational complexity, which has a direct impact on power consumption. Polar codes are favored due to their relatively low-complexity encoding and decoding algorithms. However, in practice, CRC-aided list decoders are used, which substantially increases the decoding complexity. DeepPolar is a non-linear generalization of Polar codes via neural-network-based kernels on the Plotkin tree. The DeepPolar SC-decoder is a kernel-wise sequential decoder inspired by the SC algorithm. For the case of DeepPolar $(n=256,k=37,\ell=16)$, the encoder and decoder have 0.1M and 1.6M parameters, respectively, which we did not optimize in this project. Our findings serve as constructive proof that non-linear polar codes with large kernels can outperform the state-of-the-art. Further, we demonstrate the viability of effective binary codes with a large-kernel Polar structure, complete with efficient decoders. This opens avenues for further examination of these codebooks to inform the development of low-complexity solutions. Notably, the complexity of neural codes can be reduced significantly without performance degradation by complexity-aware architecture design (eg, TinyKO (Makkuva et al., 2021)). Our preliminary experiments with DeepPolar-parallel (App. C.2), which reduces the decoder parameter count by 8x, have already demonstrated improved performance at lower complexity. Further, several techniques are used in practice, like distillation, quantization, and pruning, among others, to reduce the computational overhead of a neural algorithm. This is important and interesting future work beyond the scope of this paper.
@@ 5 Related work
The application of machine learning to channel coding has been an active area of research in recent years. The majority of existing works focus on decoding existing codes, aiming to achieve better reliability and robustness, and in some cases, lower decoding complexity (Kim et al., 2018b; Nachmani et al., 2016; Dörner et al., 2017; Vasić et al., 2018; Nachmani & Wolf, 2019; Jiang et al., 2019a; Chen & Ye, 2021; Jamali et al., 2021a; Choukroun & Wolf, 2022a, b; Hebbar et al., 2022; Aharoni et al., 2023). In the context of polar codes, several neural decoders have been proposed (Cammerer et al., 2017b, a; Doan et al., 2018; Hebbar et al., 2023). In another line of work, (Ebada et al., 2019; Liao et al., 2021; Li et al., 2021; Miloslavskaya et al., 2023; Ankireddy et al., 2024) use deep learning to identify optimal polar code frozen positions without altering the design of encoder and decoder.
In contrast, jointly learning both encoders and decoders is a more challenging problem, and very few works in the literature exist. A common theme is the incorporation of principled coding-theoretic encoding and decoding structures. TurboAE (Jiang et al., 2019b), and follow-up works (Jiang et al., 2020b, a; Chahine et al., 2021; Saber et al., 2022; Chahine et al., 2022a; Wang et al., 2023), use sequential encoding and decoding along with interleaving of input bits, inspired by Turbo codes (Berrou et al., 1993). KO codes (Makkuva et al., 2021) generalize Reed-Muller encoding and Dumer decoding by replacing selected components in the Plotkin tree with neural networks. ProductAE (Jamali et al., 2021b) generalizes two-dimensional product codes to scale neural codes to larger block lengths. In a similar vein, our work generalizes the coding structures of large-kernel Polar codes by using non-linear kernels parameterized by NNs.
Deep learning-based schemes for channels with feedback is another active area of research (Kim et al., 2018a; Safavi et al., 2021; Chahine et al., 2022b; Ozfatura et al., 2022, 2023b, 2023a; Kim et al., 2023). Furthermore, breaking the conventional wisdom that neural codes are hard to interpret, (Devroye et al., 2022) derives a closed-form approximation of binarized TurboAE codes via mixed integer linear programming and other techniques. The task of analytically approximating and understanding binarized DeepPolar codes remains a promising subject for future research.
@@ 6 Conclusion
In this work, we introduce DeepPolar codes, a new class of non-linear generalizations of large-kernel Polar codes. DeepPolar codes generalize to various rates and blocklengths, and outperform the current state-of-the-art in BER.
Impact Statement. This paper presents work whose goal is to advance the field of channel coding using machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Successive Cancellation decoder
[…]
@@
Appendix B Large kernel Polar codes
One of the drawbacks of larger kernel Polar codes in practice is the lack of efficient low-complexity decoding algorithms. The SC decoding operation for larger kernel Polar codes ($\ell>2$) is similar to the original polar codes. While the operations remain the same, iteratively computing the probabilities increases exponentially with respect to $\ell$ as $O(2^{\ell})$, making it prohibitively expensive for very large kernels (Abbasi & Viterbo, 2020; Gupta et al., 2021). Iteratively, the relation can be expressed as
@@ Appendix B Large kernel Polar codes
To demonstrate the ineffectiveness of current SC decoding strategies for larger kernel Polar codes at short-to-medium blocklengths, we consider a kernel size of $\ell=16$  (Trofimiuk & Trifonov, 2021) with SC list decoding and compare it with Polar code with SC decoding for $\ell=2$ and DeepPolar code of $\ell=16$. We see from Fig. 11 that for a $(256,37)$ code, a larger kernel Polar code is orders of magnitude worse than Polar code with $\ell=2$. Notably, the DeepPolar code of $(256,37,\ell=16)$ is better than the Polar code of $(256,37,\ell=2)$ by more than $0.5$ dB, demonstrating the efficacy of DeepPolar codes for larger kernels.
DeepPolar, $\ell$=16Polar, $\ell$=2Polar, $\ell$=16SNR (dB)
Bit Error Rate
Figure 11: $n$=256,$k$=37 : Small-to-medium block length performance of a Polar code with kernel size $\ell$=16 is sub-optimal compared to Polar and DeepPolar codes
@@ C.1 DeepPolar-SC Decoder
Figure 12: DeepPolar ($n$=16,$k$$\ell$=4) : DeepPolar-SC decoder decodes each bit sequentially write more
Figure 13: Successive Cancellation Decoding of DeepPolar (16,8,$\ell=4$).

### v2

@@ 4 Main Results
4.2 Baselines.
DeepPolar extends conventional Polar encoding and decoding structures using neural networks. We evaluate its performance against state-of-the-art learning-based codes, KO codes (Makkuva et al., 2021). Similar to DeepPolar, KO codes replaces elements of Reed-Muller by neural networks to enhance error correction capability. We also compare DeepPolar to traditional Polar codes employing Arikan’s kernel and decoded via the successive cancellation (SC) algorithm. Both codes utilize the same frozen positions, determined using the Tal-Vardy method (Tal & Vardy, 2013). We further discuss the impact of frozen position selection in App. B.1. App. B.2 discusses and compares DeepPolar to recent large-kernel polar code constructions.
@@ 4.3 Results
DeepPolar codes outperform Polar codes. As highlighted in Fig. 5, DeepPolar codes achieve enhanced performance over traditional Polar codes with SC decoding and Reed-Muller codes with Dumer decoding across a broad range of SNRs in the presence of additive white Gaussian noise (AWGN), in terms of bit error rate (BER). Notably, DeepPolar (256,37,$\ell=16$) outperforms KO($8,2$) (Fig. 5a)), the state-of-the-art neural code mapping $37$ information bits to a length-$256$ codeword (Makkuva et al., 2021) in the short-to-medium blocklength regime. DeepPolar codes, owing to their Polar-like encoding framework, accommodate a much wider range of rates than KO codes that rely on the algebraic structure of RM codes. This versatility is illustrated in Fig. 5b) and Fig. 5c), where DeepPolar codes outperform Polar codes across diverse rates and SNRs (cf. KO codes and RM codes do not exist for these rates). Further, DeepPolar is robust to non-AWGN deviations, achieving gains over Polar codes on fading channels and radar noise (App. D).
@@
Effect of kernel size. Fig. 6 highlights that scaling the kernel size $\ell$ is crucial for effective training and DeepPolar outperforming baselines. In case $n$ is not an integer exponent of $\ell$ i.e., $\ell^{m}\leq n<\ell^{m+1}$ and $\ell^{m}$ divides $n$ then the kernel size at the root is $n/\ell^{m}$. An increase in kernel size expands the function space the encoder can represent, facilitating learning more robust representations. However, empirical experiments reveal that using $\ell=\sqrt{n}$ , which results in an encoding and decoding tree of depth 2, is a good heuristic. Indeed, Fig. 1 highlights significant improvements in performance upto $\ell=16,32$, whereas further scaling $\ell=64$ leads to a decline. This trend indicates a bias-variance tradeoff, suggesting the ideal kernel size strikes a balance between model complexity and its generalization capabilities.
@@ 4.4 Interpretation
Figure 7: (n=256,k=37,$\ell=16$): Unlike Polar codes, the distribution of pairwise distribution between codewords of DeepPolar shows a strong resemblance to the Gaussian codebook. DeepPolar-binary code retains a Gaussian-like distance profile. (pdfs and pmfs have been renormalized for better visualization.)
@@
Interpreting the decoder. In our study, the training loss used is binary cross entropy (BCE), which serves as a surrogate for the bit error rate (BER). However, it is important to recognize that optimizing for BER does not necessarily translate to improved block error rate (BLER), an important figure of merit in practical systems. Our observations, as indicated in Fig. 8a), show that DeepPolar underperforms in BLER compared to baseline methods (See App. F for more results). The BLER, defined as $\mathbb{P}[\hat{\boldsymbol{u}}\neq\boldsymbol{u}]$, can be expressed as the cumulative probability of bit errors occurring at each position when no errors were made in the previous ones:
$\mathbb{P}[\hat{\boldsymbol{u}}\neq\boldsymbol{u}]=\sum_{i=1}^{k}\mathbb{P}[\hat{u}_{i}\neq u_{i}|\hat{\boldsymbol{u}}_{1:i-1}=\boldsymbol{u}_{1:i-1}]$ (3)
An analysis of error distribution (Fig. 8b)) reveals notable differences between SC and DeepPolar-SC decoders despite both employing a sequential decoding approach. Due to the effect of channel polarization, most errors in the polar decoding tend to occur at bit position 0, which predominantly drives the BLER. In contrast, errors in DeepPolar decoding are more evenly distributed across various bit positions. This uniformity is a consequence of the BCE loss prioritizing BER over BLER. This trade-off leads to a marginal decrease in BLER performance. Identifying a stable loss function directly targeting BLER, is an interesting direction for future research.
@@ 4.5 Ablation studies
DeepPolar-Binary
DeepPolar(256,37)KO(8,2)Polar(256,37)DeepPolar(256,37)-binarySNR (dB)
Bit Error Rate
Figure 9: $\ell$=16) approaches the performance of KO codes
Compared to canonical coding schemes, a real-valued coding scheme like DeepPolar offers distinct advantages by integrating modulation and coding schemes. However, practical systems may have hard symbol-level power constraints, necessitating code bits to be integers $x_{i}\in\{-1,1\}$. Hence, it is beneficial to have the ability to learn a binary code while maintaining the structure of DeepPolar.
To achieve this, a trained DeepPolar model is fine-tuned with a Straight Through Estimator (STE) (Hubara et al., 2016) and a binarization module, resulting in a binary version, DeepPolar-binary, akin to the methodology in (Jiang et al., 2019b). Fig. 9 highlights that there is a loss incurred with respect to DeepPolar. This underscores the contribution of joint coding and modulation to DeepPolar’s performance. Nevertheless, DeepPolar-binary not only surpasses the performance of the traditional Polar code but also closely matches that of KO codes. Additionally, the distance profile of DeepPolar-binary (Fig. 7) indicates it preserves Gaussian codebook-like distance properties, reinforcing the potential of non-linear binary codes derived from polarization-inducing structures. This exploration opens avenues for future research, particularly in interpreting these codes to develop novel, efficient encoders and decoders leveraging non-linear polarization-based methods.
@@
DeepPolar decoder for Polar codes
DeepPolarencoder+decoderPolar encoder+decoderPolar encoder+ DeepPolar decoderSNR (dB)
Bit Error Rate
Figure 10: For decoding Polar(256,37) codes, DeepPolar decoder matches the SC decoder.
DeepPolar codes are a generalization of Polar codes, and inherits its encoding structure (Plotkin tree, frozen set), and the sequential decoding paradigm, and achieves notable gains in BER. To pinpoint the efficacy of the learned code, we fix the Polar encoder and train the DeepPolar decoder to predict the transmitted message. Fig. 10 demonstrates that the DeepPolar decoder matches the performance of SC decoding to decode Polar codes. This is expected, owing to the SC-like parameterization of the DeepPolar decoder. The gains in BER, thus, can largely be attributed to the encoder’s non-linear, binary-to-real encoding scheme.
@@ 4.6 Computational Complexity
In our study, we introduce a novel non-linear generalization of the Polar code’s encoding and decoding structures and achieve substantial improvements in reliability. Another essential aspect of evaluating an algorithm is the computational complexity, which has a direct impact on power consumption. Polar codes are favored due to their relatively low-complexity encoding and decoding algorithms. However, in practice, CRC-aided list decoders are used, which substantially increases the decoding complexity. DeepPolar is a non-linear generalization of Polar codes via neural-network-based kernels on the Plotkin tree. The DeepPolar SC-decoder is a kernel-wise sequential decoder inspired by the SC algorithm. For the case of DeepPolar $(n=256,k=37,\ell=16)$, the encoder and decoder have 0.1M and 1.6M parameters, respectively, which we did not optimize in this project. Our findings serve as constructive proof that non-linear polar codes with large kernels can outperform the state-of-the-art. Further, we demonstrate the viability of effective binary codes with a large-kernel Polar structure, complete with efficient decoders. This opens avenues for further examination of these codebooks to inform the development of low-complexity solutions. Notably, the complexity of neural codes can be reduced significantly without performance degradation by complexity-aware architecture design (eg, TinyKO (Makkuva et al., 2021)). Our preliminary experiments with DeepPolar-parallel (App. C.2), which reduces the decoder parameter count by 8×8$\times$8 ×, have already demonstrated improved performance at lower complexity. Further, several techniques are used in practice, like distillation, quantization, and pruning, among others, to reduce the computational overhead of a neural algorithm. This is important and interesting future work beyond the scope of this paper.
@@ 5 Related work
The application of machine learning to channel coding has been an active area of research in recent years. The majority of existing works focus on decoding existing codes, aiming to achieve better reliability and robustness, and in some cases, lower decoding complexity (Kim et al., 2018b; Nachmani et al., 2016; Dörner et al., 2017; Vasić et al., 2018; Nachmani & Wolf, 2019; Jiang et al., 2019a; Chen & Ye, 2021; Jamali et al., 2021a; Choukroun & Wolf, 2022a, b; Hebbar et al., 2022; Aharoni et al., 2023). In the context of polar codes, several neural decoders have been proposed (Cammerer et al., 2017b, a; Doan et al., 2018; Hebbar et al., 2023). In another line of work, (Ebada et al., 2019; Liao et al., 2021; Li et al., 2021; Miloslavskaya et al., 2023; Ankireddy et al., 2024a) use deep learning to identify optimal polar code frozen positions without altering the design of encoder and decoder.
In contrast, jointly learning both encoders and decoders is a more challenging problem, and very few works in the literature exist. A common theme is the incorporation of principled coding-theoretic encoding and decoding structures. TurboAE (Jiang et al., 2019b), and follow-up works (Jiang et al., 2020c, a; Chahine et al., 2021; Saber et al., 2022; Chahine et al., 2022a; Wang et al., 2023), use sequential encoding and decoding along with interleaving of input bits, inspired by Turbo codes (Berrou et al., 1993). KO codes (Makkuva et al., 2021) generalize Reed-Muller encoding and Dumer decoding by replacing selected components in the Plotkin tree with neural networks. ProductAE (Jamali et al., 2021b) generalizes two-dimensional product codes to scale neural codes to larger block lengths. In a similar vein, our work generalizes the coding structures of large-kernel Polar codes by using non-linear kernels parameterized by NNs.
Deep learning-based schemes for channels with feedback is another active area of research (Kim et al., 2018a; Safavi et al., 2021; Chahine et al., 2022b; Ozfatura et al., 2022, 2023b, 2023a; Kim et al., 2023; Ankireddy et al., 2024b) Furthermore, breaking the conventional wisdom that neural codes are hard to interpret, (Devroye et al., 2022) derives a closed-form approximation of binarized TurboAE codes via mixed integer linear programming and other techniques. The task of analytically approximating and understanding binarized DeepPolar codes remains a promising subject for future research.
@@ 6 Conclusion
In this work, we introduce DeepPolar codes, a new class of non-linear generalizations of large-kernel Polar codes. DeepPolar codes generalize to various rates and blocklengths, and outperform the current state-of-the-art neural codes in BER. The neural architecture mirrors the Polar encoding and decoding structures, which along with a curriculum training methodology is key to improve generalization to unseen messages.
Acknowledgement
This research is supported in part by NSF CCF 2312753, ONR N00014-21-1-2388, NSF CNS 2112471, ARO W911NF2310062, ONR N00014-21-1-2379, NSF CNS-2008824, a gift from Intel, and Samsung Research America through 6G@@@@UT center within the Wireless Networking and Communications Group (WNCG) at the University of Texas at Austin. We thank P. Trifonov for his insightful comments.
Impact Statement
This paper presents work whose goal is to advance the field of channel coding using machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Successive Cancellation decoder
$f_{minsum}(\alpha_{1},\alpha_{2})=\textit{sign}(\alpha_{1})\textit{sign}(\alpha_{2})\textit{min}(|\alpha_{1}|,|\alpha_{2}|)$ (9)
@@
Appendix B Polar codes
DeepPolar(Tal-Vardy)Polar (Tal-Vardy)DeepPolar(5G)Polar (5G)KOPolar ($\ell$=16)SNR (dB)
Bit Error Rate
Figure 13: $n$=256,$k$=37 : The choice of frozen positions impacts the performance of DeepPolar. While using 5G frozen positions improves DeepPolar, its gains over the 5G-Polar code are modest. DeepPolar is a preliminary effort towards an automated search for large-kernel codes; it matches the best performing Polar code with $\ell=16$ (Trofimiuk & Trifonov, 2021)
B.1 5G Polar codes
Polar codes are conceptualized by channel polarization, i.e., the recursive application of the kernel results in bit-channels that are either noiseless or very noisy. However, at the practical blocklengths that we consider in this paper, there exist bit channels that are partially polarized. Consequently, the selection of frozen positions affects the reliability of polar codes. Although DeepPolar does not exhibit a clear polarization effect—due to its operation in the real field and its non-recursive kernel application—the choice of frozen positions still significantly impacts the effectiveness of the learned code.
In the paper, we consider polar codes constructed via the Tal-Vardy method as our baseline. This choice is justified by the fact that DeepPolar utilizes the Tal-Vardy rate profile, ensuring a fair comparison. However, we realize that the universal reliability sequence specified in the 5G standards can be used to find a better frozen set. Fig. 13 depicts that DeepPolar trained using the 5G frozen positions shows non-trivial improvement. However, the performance gains compared to the 5G-Polar code are modest. Exploring joint optimization of rate-matching strategies (such as freezing or precoding) along with training DeepPolar presents an intriguing avenue for future research.
B.2 Large kernel Polar codes
The SC decoding operation for larger kernel Polar codes ($\ell>2$) is similar to the original polar codes. Iteratively, the decoding equation can be expressed as
@@ B.2 Large kernel Polar codes
While these operations are similar to conventional polar codes, the complexity of iteratively computing the probabilities increases exponentially with respect to $\ell$ as $O(2^{\ell})$, making it prohibitively expensive for very large kernels. However, recent work has shown methods towards efficient decoding of a class of large-kernel polar codes (Abbasi & Viterbo, 2020; Gupta et al., 2021; Trofimiuk & Trifonov, 2021).
DeepPolar generalizes the conventional polar codes, and makes two notable modifications. First, it expands the kernel size to $\sqrt{n}$. Second, it introduces neural networks to parameterize the encoding and decoding functions for each kernel, which are then learned through training. The exhaustive search across large kernel spaces to improve upon Arikan’s original construction is infeasible. Nonetheless, recent advancements have been made by limiting this space and employing more efficient decoding methods (Trifonov, 2023). DeepPolar represents an initial attempt to automate the search for effective large-kernel polar codes.
Here, we consider an optimized construction of a kernel of size $\ell=16$  (Trofimiuk & Trifonov, 2021), with SC decoding and compare it with DeepPolar code of $\ell=16$. Fig. 13 shows that DeepPolar marginally falls short of this optimized construction. An exciting future direction would be to directly augment these codes using our methodology, akin to how KO codes augmented the Reed Muller code structure.
@@ C.1 DeepPolar-SC Decoder
Figure 14: DeepPolar ($n$=16,$k$$\ell$=4) : DeepPolar-SC decoder decodes each bit sequentially - each edge operation in the decoding tree is parameterized by a neural network $f_{d,b}$